OpenStudy (anonymous):

Hello, what is 2/3 x 2/3 x 2/3 ? Thanks in advance!

OpenStudy (anonymous):

(2/3)^3=(2*2*2)/(3*3*3)=(4*2)/(9*3)=8/27

OpenStudy (anonymous):

Thanks alot!!

OpenStudy (anonymous):

Just multiply the numerators and put them over the product of the denominators.

OpenStudy (anonymous):

thanks polpak :D